OpenStudy (anonymous):

6(cospi/6+isinpi/6) / 2(cospi/3 + isin pi/3)

OpenStudy (anonymous):

\[\frac{6 e^{\frac{i \pi }{6}}}{2 e^{\frac{i \pi }{3}}}=3 e^{-\frac{1}{6} (i \pi )}=3 \left(\cos \left(\frac{\pi }{6}\right)-i \sin \left(\frac{\pi }{6}\right)\right) \]

OpenStudy (anonymous):

We used \[ r e^ {i b} =r(\cos(b) + i \sin(b)) \]

OpenStudy (anonymous):

thanks! can you help me with this one also?? (1+squr3 i(i is not squarted)/ -1 + i If you dont want to thats ok! = )